# Barbie: Text to Barbie-Style 3D Avatars

### v2

@@ 4.5 Applications
As shown in Fig. 7-(a), our approach supports the free combination of any human body, garments and accessories, just like dressing Barbie dolls, which greatly improves the playability and reusability of text-to-avatar. With the self-evolving human prior loss, we can obtain an accurate correspondence between the generated human body and the SMPL-X model, allowing us to utilize the LBS algorithm to animate the generated avatars, as shown in Fig. 7-(b).
@@ 5 Conclusion
We propose Barbie, a novel method for creating Barbie-style 3D avatars which are dressed in disentangled, detailed, and diverse garments/accessories. To guarantee the domain-specific fidelity of the generated human body and outfits, we suitably combine different specific expert T2I models for domain-specific knowledge. To address the negative impacts caused by the over-strong generative priors of the expert models, we propose a series of well-designed loss functions and optimization strategies to ensure geometric rationality and texture harmony of the generated results. Extensive experiments demonstrate that our approach outperforms the SOTA methods in both dressed avatar and apparel generation tasks.

### v5

@@ 4.5 Applications
As shown in Fig. 8-(a), Barbie’s fine-grained decoupling enables the free combination of any human body, garments and accessories, just like dressing Barbie dolls, which greatly improves the playability and reusability of text-to-avatar. With the SMPLX-evolving prior loss, we can obtain an accurate correspondence between the generated human body and the SMPL-X model, which allows us not only to animate the generated avatars utilizing the LBS algorithm as in Fig. 8-(b), but also to easily edit garments for different body shapes as in Fig. 8-(c).
@@ 5 Conclusion
We propose Barbie, a novel framework for creating Barbie-style 3D avatars which are dressed in disentangled, detailed, and diverse garments/accessories based on prompts. To guarantee the domain-specific fidelity of the generated human body and outfits, we suitably combine different specific expert T2I models for domain-specific knowledge. To address the negative impacts caused by the over-strong generative priors of the expert models, we propose a series of well-designed loss functions and optimization strategies to ensure geometric rationality and texture harmony of the generated results. Extensive experiments demonstrate that our approach not only outperforms SOTA methods in both dressed avatar and apparel generation tasks, but also facilitates seamless composition and editing of outfits.